OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3?

OpenStudy (godlovesme):

Point-slope formula is y-y1= m(x-,x1

OpenStudy (godlovesme):

Y1 is ur given y coordinate while x1 is ir x and m is slope can u try :)

OpenStudy (anonymous):

I don't get that, its kinda confusing

OpenStudy (anonymous):

y-4?

OpenStudy (mathstudent55):

@Godlovesme meant to write: \(y - y_1 = m(x - x_1)\) Replace \(y_1\) with the y of the point, 4. Replace \(x_1\) with the x of the point, 1. Replace \(m\), the slope, with the given slope, -3

OpenStudy (godlovesme):

yeah y-y1= m(x-x1) y1--> y coordinate x1 --> x coordinate m--> slope sorry for the confusion =0

OpenStudy (anonymous):

lol its ok, im writing it out now

OpenStudy (godlovesme):

;D

OpenStudy (anonymous):

so what do i do with y-4=-3(x-1)?

OpenStudy (godlovesme):

that's it :) wait r u asked to write it in slope-intercept or standard form?

OpenStudy (anonymous):

point slope form

OpenStudy (godlovesme):

then that's all

OpenStudy (anonymous):

ooh

OpenStudy (mathstudent55):

You are correct. Good job!